# TropNNC: Structured Neural Network Compression Using Tropical Geometry

### v1

@@ Abstract
We present TropNNC, a structured pruning framework for compressing neural networks with linear and convolutional layers and ReLU activations. Our approximation is based on a geometrical approach to machine/deep learning, using tropical geometry and extending the work of Misiakos et al. (2022). We use the Hausdorff distance of zonotopes in its standard continuous form to achieve a tighter approximation bound for tropical polynomials compared to Misiakos et al. (2022). This enhancement allows for superior functional approximations of neural networks, leading to a more effective compression algorithm. Our method is significantly easier to implement compared to other frameworks, and does not depend on the availability of training data samples. We validate our framework through extensive empirical evaluations on the MNIST, CIFAR, and ImageNet datasets. Our results demonstrate that TropNNC achieves performance on par with the state-of-the-art method ThiNet, even surpassing it in compressing linear layers, and to the best of our knowledge, it is the first method that achieves this using tropical geometry.
@@ 1 Introduction
Tropical mathematics encompasses a wide range of applications, which include operations research and scheduling (Cuninghame-Green, 1979), control and optimization (Baccelli et al., 2001; Cohen et al., 2004; Butkovič, 2010; Akian et al., 2012), mathematical physics (Litvinov et al., 2001), speech recognition (Mohri et al., 2002), polyhedral geometry (Gaubert and Katz, 2011), tropical geometry (Maclagan and Sturmfels, 2021), dynamical systems on weighted lattices (Maragos, 2017), finite state transducers (Theodosis and Maragos, 2018), convex regression (Maragos and Theodosis, 2020), and machine learning (Maragos et al., 2021; Gärtner and Jaggi, 2008). Recently, tropical geometry has been applied to the theoretical study of neural networks. For example, Zhang et al. (2018) demonstrated the equivalence of ReLU-activated neural networks with tropical rational mappings, and used zonotopes to compute a bound on the number of linear regions of networks, equal to the one in (Montufar et al., 2014). Other works, like those of Charisopoulos and Maragos (2018); Alfarra et al. (2022); Smyrnis and Maragos (2020), have used tropical geometry to compute bounds on the number of linear regions in neural networks and to develop methods for pruning and compressing these networks.
@@
  • Obtaining a tighter bound for the functional approximation of tropical rational mappings than that presented in (Misiakos et al., 2022) using the standard continuous Hausdorff distance.
@@ 2 Related Work
Our framework is a structured pruning method that extends the approach of Misiakos et al. (2022). It prunes networks layer-by-layer, aiming to minimize the error introduced to the input of the next layer. In that sense, it is similar to the ThiNet framework. However, our algorithm is simpler to implement and does not require a training dataset. Our framework leverages clustering and can be naturally extended to non-uniform pruning, drawing parallels to the CUP framework. In fact, it improves upon steps 1 and 3 of CUP. For step 1, we use similar clustering vectors (referred to as filter features by CUP authors Duggal et al. (2021)) for linear layers, and introduce a new method for constructing clustering vectors in convolutional layers. Our main contribution lies in step 3, where we employ tropical geometry to select a better representative for each cluster, rather than simply choosing the neuron or channel with the most energy. While it is tempting to compare our framework with those of NISP (Yu et al., 2018) and HRank (Lin et al., 2020) given that they are both structured pruning methods, it is important to recognize the distinctions between them and ours. NISP and HRank are advanced frameworks specifically designed to address the limitations of layer-by-layer pruning approaches, such as those employed by ThiNet and our own method. In contrast, our framework leverages a rich theoretical foundation and does not rely on training data samples, offering a unique advantage in certain contexts. We believe our results are promising and present significant advantages. Our work highlights the potential of tropical geometrical techniques and paves the way for further research that could enhance their competitiveness with current state-of-the-art methods.
@@ 3.1 Background on Tropical Geometry
3.1.1 Tropical Polynomials and Rational Mappings
@@
where $[n]=\{1,...,n\}$. Here, $n$ represents the rank of the tropical polynomial. Each monomial term $\{\mathbf{a}_{i}^{T}\mathbf{x}+b_{i}\}$ of the polynomial has an exponent or slope $\mathbf{a}_{i}\in\mathbb{R}^{d}$ and a coefficient or bias $b_{i}\in\mathbb{R}$. Each monomial term corresponds to a plane in $\mathbb{R}^{d+1}$. Consequently, tropical polynomials are piecewise linear convex functions. Specifically, every tropical polynomial is a continuous piecewise linear convex function, and every continuous piecewise linear convex function can be expressed (though not uniquely) as a tropical polynomial (Maclagan and Sturmfels, 2021). The set of tropical polynomials in $\mathbf{x}$ defines the semiring $\mathbb{R}_{\mathrm{max}}[\mathbf{x}]$. Figures 1(a), 1(b) illustrate examples of tropical polynomials in one and multiple variables, respectively.
Tropical rational mappings extend the concept of rational polynomial functions to tropical algebra. A tropical rational mapping or function is defined as the tropical multiplication of a tropical polynomial $p$ by the tropical multiplicative inverse $q^{-1}$ of another tropical polynomial $q$. In conventional arithmetic, this operation corresponds to the difference of two tropical polynomials $p$ and $q$:
@@
Tropical rational mappings correspond to general piecewise linear functions. Specifically, every tropical rational mapping is a continuous piecewise linear function, and every continuous piecewise linear function can be expressed (though not uniquely) as a tropical rational mapping. Figures 1(c), 1(d) provide examples of tropical rational mappings.
@@
Figure 1: (a) depicts a single-variate tropical polynomial, (b) depicts a multi-variate tropical polynomial, (c) depicts a single-variate tropical rational mapping, (d) depicts a multi-variate tropical rational mapping
@@ 3.1.2 Newton Polytopes
Let $p\in\mathbb{R}_{\mathrm{max}}[\mathbf{x}]$ a tropical polynomial in $d$ variables with extended Newton polytope $P=\mathrm{ENewt}(p)$. If $(\mathbf{a}^{T},b)$ lies below the upper envelope of $P$, then $\forall\mathbf{x}\in\mathbb{R}^{d},\mathbf{a}^{T}\mathbf{x}+b\leq p(\mathbf{x})$. The inequality is strict if $(\mathbf{a}_{i}^{T},b_{i})$ lies strictly below the upper envelope.
@@
which correspond to the vertices of the upper envelope of $\mathrm{ENewt}(f\vee g)$.
@@ 3.1.3 Zonotopes
where $I\subset[n]$. It can be shown that zonotopes have up to exponential many verticesIn fact, for generators in general positions it has exactly $2\sum_{j=0}^{d-1}nCr(n-1,j)$ vertices (Gritzmann and Sturmfels, 1993).
@@ 3.2 Neural Networks with Piecewise Linear Activations
In the above expression, $|c_{ji}|f_{i}$ are tropical polynomials. Thus, $p_{j}$ and $q_{j}$ are tropical polynomials formed by the addition of tropical polynomials. Consequently, $v_{j}$ is a tropical rational mapping. We call $p_{j}$ the positive polynomial and $q_{j}$ the negative polynomial of $v_{j}$. This result can be extended to deeper networks, as suggested by the following proposition:
@@
A ReLU-activated deep neural network $\mathbf{F}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{m}$ is a tropical rational mapping.
@@ 4 Approximation based on Hausdorff distance
In this section, we present our improved theorem, which uses the Hausdorff distance in its standard continuous form to bound the error between the original and approximate neural networks. Misiakos et al. (2022) derived a bound for the error in approximating the tropical polynomials that represent a neural network. The motivation for geometrically bounding the error of these polynomials stems from Theorem 5. This theorem indicates that two polynomials with the same extended Newton polytopes are functionally equivalent. Consequently, it is expected that two tropical polynomials with approximately equal extended Newton polytopes will attain similar values.
@@
Since $V_{P}\subset P,V_{\tilde{P}}\subset\tilde{P}$, we have that $\mathrm{dist}(\mathbf{u},\tilde{P})\leq\mathrm{dist}(\mathbf{u},V_{\tilde{P}}),\mathrm{dist}(P,\mathbf{v})\leq\mathrm{dist}(V_{P},\mathbf{v})$. The equality in the aforementioned inequalities is achieved only in special cases and thus the bound $H(P,\tilde{P})$ we provide is in general tighter than $DH(P,\tilde{P})$ from Misiakos et al. (2022). We demonstrate one example where the two formulations give equal distances and one where the standard Hausdorff distance is strictly smaller than the discrete one.
@@
Notice that if our goal were to approximate the polytope $P$ by the polytope $\tilde{P}$, then according to the discrete Hausdorff distance, taking the mean is the optimal strategy and gives distance $\sqrt{3}/2$ (example 2), better than $1$ (example 3). On the other hand, according to the Hausdorff distance in its standard form, taking the sum is the optimal strategy and gives distance $1/2$ (example 2), better than $\sqrt{3}/2$ (example 3). This implies that an algorithm based on the Hausdorff distance would imply a better approximation of $\tilde{P}$ in the above example.
We derive the following result, which differs from (Misiakos et al., 2022) by using the standard form of the Hausdorff distance and thus providing a tighter bound. For the proof we apply triangle inequality and then repeatedly Proposition 10 to each output of the network. The full proofs of Proposition 10 and Theorem 11 are provided in Appendix A.3.
@@ 5.1 Single Output
First, we focus on single output networks, as the one depicted in Figure 3, with $m=1$. The network has weight matrices $\mathbf{A}\in\mathbb{R}^{n\times d}$, $\mathbf{C}\in\mathbb{R}^{1\times n}$, and a single output $v$ with corresponding positive and negative zonotopes $P$ and $Q$. We aim to approximate this network with a smaller one that has fewer neurons in the hidden layer, say $K<n$, resulting in a new hidden layer $\tilde{\mathbf{f}}=(\tilde{f}_{1},\ldots,\tilde{f}_{K})$ with new weight matrices $\widetilde{\mathbf{A}}$ and $\widetilde{\mathbf{C}}$. Our goal is to ensure that $\tilde{v}(\mathbf{x})\approx v(\mathbf{x})$ for all $\mathbf{x}\in B$. According to Theorem 11, it is sufficient to choose the new weight matrices $\widetilde{\mathbf{A}}$ and $\widetilde{\mathbf{C}}$ such that $\widetilde{P}\approx P$ and $\widetilde{Q}\approx Q$, where the approximation relation is in terms of the Hausdorff distance between the zonotopes.
@@ 5.1.1 Algorithm for single output network
Figure 5: Example execution of Zonotope K-means and Improved Zonotope K-means
@@ 5.2.1 Non-iterative Algorithm for multi-output network
There are several variations of Algorithm 2. For instance, one could use the sum instead of the mean for the input weights. However, in our experiments, this approach yielded worse results. Another variation involves clustering neurons based on cosine similarity rather than the L2 distance of the vectors, recognizing that neurons with parallel input weights can be clustered together regardless of their output weights. Additionally, one could normalize the vectors based on their input weights by shifting the norm to the output weights before applying K-means. Despite testing these methods, none provided significant advantages. Therefore, we simply recommend Algorithm 2.
@@ 5.2.2 Iterative Algorithm for multi-output network
$\sum_{j=1}^{m}\left[\tilde{C}_{jk}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})-\sum_{i\in I_{k}}C_{ji}(\mathbf{a}_{i}^{T},b_{i})\right]^{2},$
@@
6 Discussion and Limitations
@@
As for other limitations, we should note that our algorithm is designed to work strictly on layers as described above. Namely, convolutional or linear layers with ReLU and maxout or average pooling layers in between. Layers that apply computations, like batch normalization layers or the loops of residual blocks of ResNet networks fundamentally change the theory for the analysis of the network. For example, similary to ThiNet, our methods can not be applied for the compression of the terminal convolutional layer of residual blocks.
Non-uniform pruning. The techniques we have presented apply a uniform pruning ratio to all layers of the network. However, they can easily be extended to non-uniform pruning if instead of K-means we use hierarchical clustering with a global threshold parameter, like in the CUP framework (Duggal et al., 2021). In fact, when used with hierarchical clustering, TropNNC differs from CUP in step 1, where we choose a different approach to build the cluster vectors (or filter features as Duggal et al. (2021) call them) of a convolutional layer, and more importantly in step 3, where we choose a different cluster representative based on tropical geometry.

### v2

@@ Abstract
We present TropNNC, a framework for compressing neural networks with linear and convolutional layers and ReLU activations. TropNNC is a structured compression framework based on a geometrical approach to machine/deep learning, using tropical geometry and extending the work of Misiakos et al. (2022). We use the Hausdorff distance of zonotopes in its standard continuous form to achieve a tighter approximation bound for tropical polynomials compared to previous work. This enhancement leads to the development of an effective compression algorithm that achieves superior functional approximations of neural networks. Our method is significantly easier to implement compared to other frameworks, and does not depend on the availability of training data samples. We validate our framework through extensive empirical evaluations on the MNIST, CIFAR, and ImageNet datasets. Our results demonstrate that TropNNC achieves performance on par with state-of-the-art methods like ThiNet (even surpassing it in compressing linear layers) and CUP. To the best of our knowledge, it is the first method that achieves this using tropical geometry.
@@ 1 Introduction
Tropical mathematics encompasses a wide range of applications, which include operations research and scheduling (Cuninghame-Green, 1979), control and optimization (Baccelli et al., 2001; Cohen et al., 2004; Butkovič, 2010; Akian et al., 2012), mathematical physics (Litvinov et al., 2001), speech recognition (Mohri et al., 2002), polyhedral geometry (Gaubert and Katz, 2011), tropical geometry (Maclagan and Sturmfels, 2021), dynamical systems on weighted lattices (Maragos, 2017), sparsity theory on weighted lattices (Tsilivis et al., 2022), matrix factorization (Kordonis et al., 2024), finite state transducers (Theodosis and Maragos, 2018), convex regression (Maragos and Theodosis, 2020), and machine learning (Maragos et al., 2021; Gärtner and Jaggi, 2008). Recently, tropical geometry has been applied to the theoretical study of neural networks. For example, Zhang et al. (2018) demonstrated the equivalence of ReLU-activated neural networks with tropical rational mappings, and used zonotopes to compute a bound on the number of linear regions of networks, equal to the one in (Montufar et al., 2014). Other works, like those of Charisopoulos and Maragos (2018); Alfarra et al. (2023); Smyrnis and Maragos (2020), have used tropical geometry to compute bounds on the number of linear regions in neural networks and to develop methods for pruning and compressing these networks.
@@
  • Obtaining a tighter bound for the functional approximation of tropical polynomials than that presented in (Misiakos et al., 2022) using the standard, continuous Hausdorff distance.
@@ 2 Related Work
Our framework is a structured pruning method that extends the approach of Misiakos et al. (2022). It prunes networks layer-by-layer, aiming to minimize the error introduced to the input of the next layer. In that sense, it is similar to the ThiNet framework. However, our algorithm is simpler to implement and does not require a training dataset. Our framework leverages clustering and can be naturally extended to non-uniform pruning, drawing parallels to the CUP framework. In fact, it improves upon steps 1, 2, and 3 of CUP. For step 1, we use similar clustering vectors (referred to as filter features by CUP authors Duggal et al. (2021)) for linear layers, and introduce a new method for constructing clustering vectors in convolutional layers. For step 2, we introduce two approaches for choosing the distance threshold of the hierarchical clustering. Our main contribution lies in step 3, where we employ tropical geometry to select a better representative for each cluster, rather than simply choosing the neuron or channel with the most energy. While it is tempting to compare our framework with those of NISP (Yu et al., 2018) and HRank (Lin et al., 2020) given that they are both structured pruning methods, it is important to recognize the distinctions between them and ours. NISP and HRank are advanced frameworks specifically designed to address the limitations of layer-by-layer pruning approaches, such as those employed by ThiNet and our own method. In contrast, our framework leverages a rich theoretical foundation and does not rely on training data samples, offering a unique advantage in certain contexts. We believe our results are promising and present significant advantages. Our work highlights the potential of tropical geometrical techniques and paves the way for further research that could enhance their competitiveness with current state-of-the-art methods.
@@ 3.1 Background on Tropical Geometry
3.1.1 Tropical Polynomials and Rational Functions
@@
where $[n]:=\{1,...,n\}$. Here, $n$ represents the rank of the tropical polynomial. Each monomial term $\{\mathbf{a}_{i}^{T}\mathbf{x}+b_{i}\}$ of the polynomial has an exponent or slope $\mathbf{a}_{i}\in\mathbb{R}^{d}$ and a coefficient or bias $b_{i}\in\mathbb{R}$. Each monomial term corresponds to a plane in $\mathbb{R}^{d+1}$. Consequently, tropical polynomials are piecewise linear convex functions. Specifically, every tropical polynomial is a continuous piecewise linear convex function, and every continuous piecewise linear convex function can be expressed (though not uniquely) as a tropical polynomial (Maclagan and Sturmfels, 2021). The set of tropical polynomials in $\mathbf{x}$ defines the semiring $\mathbb{R}_{\mathrm{max}}[\mathbf{x}]$. Figures 1(a), 1(b) illustrate examples of tropical polynomials in one and multiple variables, respectively.
Tropical rational functions extend the concept of rational polynomial functions to tropical algebra. A tropical rational function is defined as the tropical multiplication of a tropical polynomial $p$ by the tropical multiplicative inverse $q^{-1}$ of another tropical polynomial $q$. In conventional arithmetic, this operation corresponds to the difference of two tropical polynomials $p$ and $q$:
@@
Tropical rational functions correspond to general piecewise linear functions. Specifically, every tropical rational function is a continuous piecewise linear function, and every continuous piecewise linear function can be expressed (though not uniquely) as a tropical rational function. Figures 1(c), 1(d) provide examples of tropical rational functions.
@@
Figure 1: (a) depicts a single-variate tropical polynomial, (b) depicts a multi-variate tropical polynomial, (c) depicts a single-variate tropical rational function, (d) depicts a multi-variate tropical rational function
@@ 3.1.2 Newton Polytopes
Let $p\in\mathbb{R}_{\mathrm{max}}[\mathbf{x}]$ a tropical polynomial in $d$ variables with extended Newton polytope $P=\mathrm{ENewt}(p)$. If $(\mathbf{a}^{T},b)$ lies below the upper envelope of $P$, then $\forall\mathbf{x}\in\mathbb{R}^{d},\mathbf{a}^{T}\mathbf{x}+b\leq p(\mathbf{x})$. The inequality is strict if $(\mathbf{a}^{T},b)$ lies strictly below the upper envelope.
@@
which corresponds to the vertices of the upper envelope of $\mathrm{ENewt}(f\vee g)$.
@@ 3.1.3 Zonotopes
where $I\subseteq[n]$. It can be shown that zonotopes have up to exponential many verticesIn fact, for generators in general positions they have exactly $2\sum_{j=0}^{d-1}nCr(n-1,j)$ vertices (Gritzmann and Sturmfels, 1993).
@@ 3.2 Neural Networks with Piecewise Linear Activations
In the above expression, $|c_{ji}|f_{i}$ are tropical polynomials. Thus, $p_{j}$ and $q_{j}$ are tropical polynomials formed by the addition of tropical polynomials. Consequently, $v_{j}$ is a tropical rational function. We call $p_{j}$ the positive polynomial and $q_{j}$ the negative polynomial of $v_{j}$. This result can be extended to deeper networks, as suggested by the following proposition:
@@
A ReLU-activated deep neural network $\mathbf{F}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{m}$ is a tropical rational mapping (vector whose elements are tropical rational functions).
@@ 4 Approximation based on Hausdorff distance
In this section, we present our improved theorem, which uses the Hausdorff distance in its standard continuous form to bound the error between the original and approximate neural networks. Misiakos et al. (2022) derived a bound for the error of approximating the tropical polynomials that represent a neural network. The motivation for geometrically bounding the error of these polynomials stems from Theorem 5. This theorem indicates that two polynomials with the same extended Newton polytopes are functionally equivalent. Consequently, it is expected that two tropical polynomials with approximately equal extended Newton polytopes will attain similar values.
@@
Since $V_{P}\subseteq P,V_{\tilde{P}}\subseteq\tilde{P}$, we have that $\mathrm{dist}(\mathbf{u},\tilde{P})\leq\mathrm{dist}(\mathbf{u},V_{\tilde{P}}),\mathrm{dist}(P,\mathbf{v})\leq\mathrm{dist}(V_{P},\mathbf{v})$. The equality in the aforementioned inequalities is achieved only in special cases and thus the bound $H(P,\tilde{P})$ we provide is in general tighter than $DH(P,\tilde{P})$ from Misiakos et al. (2022). We demonstrate one example where the two formulations give equal distances and one where the standard Hausdorff distance is strictly smaller than the discrete one.
@@
Notice that if our goal were to approximate the polytope $P$ by the polytope $\tilde{P}$, then according to the discrete Hausdorff distance, taking the mean is the optimal strategy and gives distance $\sqrt{3}/2$ (example 2), better than $1$ (example 3). On the other hand, according to the Hausdorff distance in its standard form, taking the sum is the optimal strategy and gives distance $1/2$ (example 2), better than $\sqrt{3}/2$ (example 3). This implies that an algorithm based on the Hausdorff distance would imply a better approximation of $P$ in the above example.
We derive the following result, which differs from (Misiakos et al., 2022) by using the standard form of the Hausdorff distance and thus providing a tighter bound. For the proof we apply the triangle inequality and then repeatedly Proposition 10 to each output of the network. The full proofs of Proposition 10 and Theorem 11 are provided in Appendix A.3.
@@ 5.1 Single Output
First, we focus on single output networks, as the one depicted in Figure 3, with $m=1$. The network has weight matrices $\mathbf{A}\in\mathbb{R}^{n\times(d+1)}$, $\mathbf{C}\in\mathbb{R}^{1\times n}$, and a single output $v$ with corresponding positive and negative zonotopes $P$ and $Q$. We aim to approximate this network with a smaller one that has fewer neurons in the hidden layer, say $K<n$, resulting in a new hidden layer $\tilde{\mathbf{f}}=(\tilde{f}_{1},\ldots,\tilde{f}_{K})$ with new weight matrices $\widetilde{\mathbf{A}}$ and $\widetilde{\mathbf{C}}$. Our goal is to ensure that $\tilde{v}(\mathbf{x})\approx v(\mathbf{x})$ for all $\mathbf{x}\in B$. According to Theorem 11, it is sufficient to choose the new weight matrices $\widetilde{\mathbf{A}}$ and $\widetilde{\mathbf{C}}$ such that $\widetilde{P}\approx P$ and $\widetilde{Q}\approx Q$, where the approximation relation is in terms of the Hausdorff distance between the zonotopes.
@@ 5.1.1 Algorithm for single output network
Figure 5: Example execution of Zonotope K-means and TropNNC
@@ 5.2.1 Non-iterative Algorithm for multi-output network
There are several variations of Algorithm 2. For instance, one could use the sum instead of the mean for the input weights. However, in our experiments, this approach yielded worse results. Another variation involves clustering neurons based on cosine similarity rather than the $L^{2}$ distance of the vectors, recognizing that neurons with parallel input weights can be clustered together regardless of their output weights. Additionally, one could normalize the vectors based on their input weights by shifting the norm to the output weights before applying K-means. Despite testing these methods, none provided significant advantages. Therefore, we simply recommend Algorithm 2.
@@ 5.2.2 Iterative Algorithm for multi-output network
$\sum_{j=1}^{m}\left\|\tilde{C}_{jk}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})-\sum_{i\in I_{k}}C_{ji}(\mathbf{a}_{i}^{T},b_{i})\right\|^{2},$
@@
6 Extensions, Limitations, and Applications of the Method
@@
Using these ideas, we can fuse a batch normalization layer with its preceding convolutional or linear layer. This enables two approaches for applying our algorithm to networks with batch normalization. The first approach involves fusing all layers of the network beforehand, resulting in a fully fused network without batch normalization. The second approach performs fusion during the compression process, applying it layer-by-layer as each layer is compressed. For the second approach, we made a notable observation during our experiments, although the reasoning behind it remains unclear: using the pre-fusion weights/kernels of the layers for clustering vectors consistently led to significantly better results.
As for other limitations, we should note that our algorithm is designed to work strictly on layers as described above. Namely, convolutional or linear layers with ReLU and maxout or average pooling layers in between. Layers that apply computations, like batch normalization layers or the loops of residual blocks of ResNet networks fundamentally change the theory for the analysis of the network. For example, similarly to ThiNet, our methods can not be applied for the compression of the terminal convolutional layer of residual blocks.
Non-uniform compression. The techniques we have presented apply a uniform pruning ratio to all layers of the network. However, they can easily be extended to non-uniform pruning if instead of K-means we use hierarchical clustering with a global threshold parameter, like in the CUP framework (Duggal et al., 2021). In fact, when used with hierarchical clustering, TropNNC differs from CUP in step 1, where we choose a different approach to build the clustering vectors (or filter features as Duggal et al. (2021) call them) of a convolutional layer, and more importantly in step 3, where we choose a different cluster representative based on tropical geometry. We also propose a modification to step 2 of CUP, specifically the hierarchical clustering step using a global distance threshold. Since the clustering vectors of different layers have varying dimensions, and vectors in higher-dimensional spaces tend to be more spread out, we introduce two variants for selecting the distance threshold for each layer:
  • Variant 1: For each layer, take the distance threshold to be some global constant times the square root of the dimension of the clustering vectors of the layer.
  • Variant 2: For each layer, take the distance threshold to be some global constant times the mean of the norms of the clustering vectors of the layer.
As demonstrated in our experiments, the first variant excels when minimizing the number of parameters, while the second variant is more effective when our objective is to minimize floating-point operation demands.